x = 3(5+2x): x = 3(5+2x)
Answer by jim_thompson5910(9401)

x=3(5+2x) Start with the given equation.


x=15+6x Distribute.


x-6x=15 Subtract 6x from both sides.


-5x=15 Combine like terms on the left side.


x=(15)/(-5) Divide both sides by -5 to isolate x.


x=-3 Reduce.


----------------------------------------------------------------------

Answer:

So the answer is x=-3
Question 152540: x = 3(5+2x): x = 3(5+2x)
Answer by checkley77(3654)

x=3(5+2x)
x=15+6x
x-6x=15
-5x=15
x=15/-5
x=-3 answer.
proof:
-3=3(5+2*-3)
-3=3(5-6)
-3=3*-1
-3=-3